# Efficient and operational quantifier of divisibility in terms of channel discrimination

### v1

@@ IV Divisibility as semi-definite program
It is exactly the previous corollary that allows us to reformulate the question of whether or not a given quantum dynamics =$\mathcal{F}=\{\mathcal{N}_{k|0}\}_{k\in[N]}$ is P or CP divisible. Focusing on the image of the quantum channels via the Jamiołkowski isomorphism, the question about the CP-divisibility of two given quantum maps, $\varrho_{2|0}$ and $\varrho_{1|0}$, can be re-cast as SDPs, as shown below.
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0}^{AC},\varrho_{1|0}^{AB}$
$\text{Find}\leavevmode\nobreak\ \varrho_{2|1}^{BC}$
s.t. (15)
$\qquad\varrho_{2|0}=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\varrho_{2|1}^{BC})^{T_{B}}(\varrho_{1|0}^{AB}\otimes\mathds{1}_{C})\right]$
$\qquad\varrho_{2|1}^{BC}\geq 0$
$\qquad\mbox{Tr}_{C}\left[\varrho_{2|1}^{BC}\right]=\mathds{1}_{B}.$
@@
An operationally meaningful measure for the distance is the diamond norm Watrous (2018, 2012), which can be cast as an SDP and is directly related to the task of optimally distinguishing two channels when entangled resources are provided Pirandola et al. (2019); Brandão et al. (2015) (see Appendix C for details). We represent the diamond norm by $\|\bullet\|_{\diamond}$.
As shown below, the only modification from the SDP (IV) is the replacement of the equality requirement $\Lambda_{2|0}=\Lambda_{2|1}^{BC}\circ\Lambda_{1|0}$, which can only be satisfied if $\Lambda_{2|0}$ is divisible with respect to $\Lambda_{1|0}$, with a minimization of the distance between the two sides of the equation. This replacement introduces a "soft" constraint that penalizes deviations from the equality constraint but keeps the problem feasible even when the maps are not exactly divisible. This formulation of the problem, stated below, is the one we employed in the results obtained in the next sections:
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0}^{AC},\varrho_{1|0}^{AB}$
$\operatorname*{Minimise}_{\varrho_{2|1}^{BC}}\|\varrho_{2|0}^{AC}-\varrho_{2|1}^{BC}\circ\varrho^{AB}_{1|0}\|_{\diamond}$
s.t. (16)
$\qquad\varrho_{2|1}^{BC}\geq 0$
$\qquad\mbox{Tr}_{C}\left[\varrho_{2|1}^{BC}\right]=\mathds{1}_{B}.$
Analogously, in order to decide about the P-divisibility of two given quantum maps $\rho_{2|0}^{AC},\rho_{1|0}^{AB}$, we can reformulate the above problem, relaxing accordingly to also allow for positive (but not completely positive) linear maps.
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\varrho_{1|0}$
$\text{Find}\leavevmode\nobreak\ \varrho_{2|1}$
s.t. (17)
$\qquad\varrho_{2|0}=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\varrho_{2|1})^{T_{B}}(\varrho_{1|0}\otimes\mathds{1}_{C})\right]$
$\qquad\sum_{i\neq j}\left[\lambda_{i}(\varrho_{2|1})+\lambda_{j}(\varrho_{2|1})\right]\geq 0$
$\qquad\mbox{Tr}_{C}\left[\varrho_{2|1}\right]=\mathds{1}_{B},$
where $\lambda_{k}(\varrho_{2|1})$ is the $k-$th eigenvalue of the matrix $\varrho_{2|1}$
@@
for all separable states $\sigma$ Bengtsson and Zyczkowski (2006). Effectively, this means that if $\varrho$ is the image of a PTP but non-CPTP map, it must act as an entanglement witness, since some entangled state $\sigma$ will fail to preserve the positivity of the expression above. To turn the P-divisibility test into an SDP, therefore requires a description (or an approximate description) of the set of entanglement witnesses in terms of linear matrix equalities and inequalities.
Although a general characterization of positive maps is expected to be a hard problem, given its connection to the problem of separability, a semidefinite program characterization of maps whose domain and range are 2-dimensional Hilbert spaces is possible, due to Størmer’s theorem Majewski and Marciniak (2007); Størmer (1982); Woronowicz (1976). This means replacing the positive-semidefiniteness constraint in Eq. (IV) for $\Lambda_{2|1}$ with the decomposability relation,
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\varrho_{1|0}$
$\text{Find}\leavevmode\nobreak\ \varrho_{2|1}$
s.t. (20)
$\qquad\varrho_{2|0}=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\varrho_{2|1})^{T_{B}}(\varrho_{1|0}\otimes\mathds{1}_{C})\right]$
$\qquad\varrho_{2|1}=\sigma_{A}+\sigma_{B}$
$\qquad\sigma_{A}\geq 0$
$\qquad\sigma_{B}\mbox{ is the Jamio\l kowski image of }\Lambda_{2|1}\circ T_{1}$
$\qquad\mbox{Tr}_{C}\left[\varrho_{2|1}\right]=\mathds{1}_{B},$
where $\Lambda_{2|1}$ is CPTP and $T_{1}$ is the partial transposition on the first system on a given local basis.
@@ V Witnessing non-Markovianity in a Quantum Dynamics
Non-markovian effects are commonly associated with a backflow of information Chruściński et al. (2011). One of the most typical measures of it relies on the increase of distinguishability of two states Breuer et al. (2009). Differently, the formalism proposed here does not rely on any witness of the backflow of information and does not request any restriction on the dynamics. To show its relevance, we will apply it to two paradigmatic models.
The first example is a collisional model. This class of model has been initially proposed as a toy model to simulate any Markovian dynamics. Recently, by allowing the generation of correlations in the environment, collisional models have been proven to be a powerful tool in the studies of non-Markovian dynamics (for a review, see Ciccarello et al. (2022)). In this model the environment is described by separated particles and the system interacts with each particle at once and sequentially—see Fig. 2. The advantage of this model is that the three main ingredients that may lead to a non-Markovian evolution, namely the correlations in the environment, its internal dynamics, and the detailed interaction between the system and the environment can be separately fully controllable. Indeed, for qubits, collisional models can simulate any non-Markovian dynamics Rybár et al. (2012).
@@ V.1 Collisional Model
Two main ingredients are common in all collisional models: time is discretized, and the environment is considered to be formed by elementary subsystems. Each encounter between the system and the $j$-th particle in the environment is denoted by the unitary operator $U_{j}$, acting for a fixed duration $\tau$. The environment’s particles are typically identical and exhibit no interactions. The resulting evolution after $j$ collisions is expressed as $\rho_{j\tau|0}=\text{Tr}_{j}...\text{Tr}_{1}[U_{j}...U_{1}(\rho\otimes\omega_{env})U_{1}^{\dagger}...U_{j}^{\dagger}]$. A pictorial representation of this process is displayed in Fig. 2. This process simulates a Markovian dynamics. However, for non-Markovian dynamics, as indicated in Rybár et al. (2012); Ciccarello et al. (2013); Bernardes et al. (2014), the environment’s particles can no longer remain noninteracting; instead, they must exhibit some level of correlation.
@@
An advantage of using SDPs to quantify P- and CP-divisibility is that the best channel approximating the optimal distance is obtained as a by-product. In some cases, however, the resulting best channel has been observed to be the identity, meaning that no other channel will improve in approximating $\Lambda_{2|0}$ starting at $\Lambda_{1|0}$, $\Lambda_{1|0}$ being already the closest channel as a factor composing the former. This induces a notion of absolute non-divisibility and a possible justification for the result is introduced in Appendix D.
@@
Figure 5: Distance to divisibility for the dephasing channel. Dynamics starts at time $t_{1}$ and ends at time $t_{2}$ (with values in the range $[0,2\pi]$ and $t_{2}\geq t_{1}$). On the left panel, divisibility for a selection of initial times is shown. Notice that, as $t_{1}$ increases, the sooner non-divisibility starts, with extended regions of exact divisibility occurring only for $t_{1}\leq\pi$. Compared with figure 4, it can be seen that divisibility can only be attained when $p(t_{2})>p(t_{1})$.
@@ V.2 Dephasing Dynamics
The dephasing channel is a completely quantum noise channel, where quantum information is lost without the loss of energy. It models fiber-optic communication channels Erhard et al. (2020) and superconducting circuits Brito et al. (2008), impacting quantum communication and quantum computation.
@@
with $p(t)=e^{4}\,(1-e^{-2(1-\cos(t))})/(e^{4}-1)$ (see Fig. 4). The corresponding master equation is $\dot{\rho}=\frac{\gamma(t)}{2}(Z\rho Z-\rho)$ with $\gamma(t)\leq 0$ for $t\geq\pi$ and consequently non-Markovian Chruściński and Wudarski (2013). Another interesting feature of this model is that for $p=1$ the map is non-invertible; this is a regime where many different Markovian witnesses are not well defined De Santis et al. (2019, 2020); Kołodyński et al. (2020). The time parameter here is again continuous, but note that we will always be interested in three different time steps: the initial time $t_{0}$, the intermediate time $t_{1}$, and the final time $t_{2}$.

### v2

@@ IV Divisibility as semi-definite program
It is exactly the previous corollary that allows us to reformulate the question of whether or not a given quantum dynamics $\mathcal{F}=\{\mathcal{N}_{k|0}\}_{k\in[N]}$ is P or CP divisible. Focusing on the image of the quantum channels via the Jamiołkowski isomorphism, the question about the CP-divisibility of two given quantum maps, $\varrho_{2|0}$ and $\varrho_{1|0}$, can be re-cast as SDPs, as shown below.
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\varrho_{1|0}$ $\text{Find}\leavevmode\nobreak\ \varrho_{2|1}$ s.t. (15) $\qquad\varrho_{2|0}=\mbox{Tr}_{B}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{2|1})^{T_{\mathcal{H}_{1}}}(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right]$ $\qquad\varrho_{2|1}\geq 0$ $\qquad\mbox{Tr}_{\mathcal{H}_{2}}\left[\varrho_{2|1}\right]=\mathds{1}_{\mathcal{H}_{1}}.$
@@
An operationally meaningful measure for the distance is the diamond norm [44, 45], which can be cast as an SDP and is directly related to the task of optimally distinguishing two channels when entangled resources are provided [30, 31] (see Appendix C for details). We represent the diamond norm by $\|\bullet\|_{\diamond}$. An advantage of reformulating the SDP in terms of distance to divisibility is that it admits an equivalent dual form which provides, as subproducts of the optimization, a pair of operators $(W,\Lambda)$ that work together as a witness for CP-divisibility, for a fixed $\varrho_{1|0}$. They give rise to the necessary condition $\mbox{Tr}[W\,\varrho_{2|0}]+\mbox{Tr}[\Lambda]\leq 0$ for divisibility, although limited to the given pair of maps $\varrho_{2|0}$ and $\varrho_{1|0}$. Details of the dual form and witness are presented also in Appendix C.
As shown below, the only modification from the SDP (IV) is the replacement of the equality requirement $\Lambda_{2|0}=\Lambda_{2|1}\circ\Lambda_{1|0}$, which can only be satisfied if $\Lambda_{2|0}$ is (P- or CP-)divisible with respect to $\Lambda_{1|0}$, with a minimization of the distance between the two sides of the equation. This replacement introduces a “softer” constraint that penalizes deviations from the equality constraint but keeps the problem feasible even when the maps are not exactly divisible. This formulation of the problem, stated below, is the one we employed in the results obtained in the next sections:
CP-Divisibility with Diamond-Norm
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\,\varrho_{1|0}$ $\operatorname*{Minimise}_{\varrho_{2|1}}\|\mathcal{J}^{-1}(\varrho_{2|0})-\mathcal{J}^{-1}(\varrho_{2|1})\circ\mathcal{J}^{-1}(\varrho_{1|0})\|_{\diamond}$ s.t. (16) $\qquad\varrho_{2|1}\geq 0$ $\qquad\mbox{Tr}_{\mathcal{H}_{2}}\left[\varrho_{2|1}\right]=\mathds{1}_{\mathcal{H}_{1}}.$ where $\mathcal{J}^{-1}(\varrho_{k|0})$ is the map associated with $\varrho_{k|0}$ via the (Choi)-Jamiołkowski isomorphism.
Analogously, in order to decide about the P-divisibility of two given quantum maps $\rho_{2|0},\rho_{1|0}$, we can reformulate the above problem, relaxing accordingly to also allow for positive (but not completely positive) linear maps.
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\,\varrho_{1|0}$ $\text{Find}\leavevmode\nobreak\ \varrho_{2|1}$ s.t. (17) $\qquad\varrho_{2|0}=\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{2|1})^{T_{\mathcal{H}_{1}}}(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right]$ $\qquad\sum_{i\neq j}\left[\lambda_{i}(\varrho_{2|1})+\lambda_{j}(\varrho_{2|1})\right]\geq 0$ $\qquad\mbox{Tr}_{\mathcal{H}_{2}}\left[\varrho_{2|1}\right]=\mathds{1}_{\mathcal{H}_{1}},$ where $\lambda_{k}(\varrho_{2|1})$ is the $k$-th eigenvalue of the matrix $\varrho_{2|1}$
@@
for all separable states $\sigma$ [46]. Effectively, this means that if $\varrho$ is the image of a PTP but non-CPTP map, it must act as an entanglement witness, since some entangled state $\sigma$ will fail to preserve the positivity of the expression above. To turn the P-divisibility test into an SDP, therefore requires a description (or an approximate description) of the set of entanglement witnesses in terms of linear matrix equalities and inequalities. The connection between the generalized robustness of entanglement, an entanglement measure that can be calculated by SDP, and divisibility has also been explored in [18].
Although a general characterization of positive maps is expected to be a hard problem, given its connection to the problem of separability, a semidefinite program characterization of maps whose domain and range are 2-dimensional Hilbert spaces is possible, due to Størmer’s theorem [47, 48, 49]. This means replacing the positive-semidefiniteness constraint in Eq. (IV) for $\Lambda_{2|1}$ with the decomposability relation,
@@
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\,\varrho_{1|0}$ $\text{Find}\leavevmode\nobreak\ \varrho_{2|1},\,\sigma_{A},\,\sigma_{B}$ s.t. (20) $\qquad\varrho_{2|0}=\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{2|1})^{T_{\mathcal{H}_{1}}}(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right]$ $\qquad\varrho_{2|1}=\sigma_{A}+\sigma_{B}$ $\qquad\sigma_{A}\geq 0$ $\qquad\sigma_{B}\mbox{ is the Jamio\l kowski image of }\Lambda_{2|1}\circ T_{1}$ $\qquad\mbox{Tr}_{\mathcal{H}_{2}}\left[\varrho_{2|1}\right]=\mathds{1}_{\mathcal{H}_{1}},$ where $\Lambda_{2|1}$ is CPTP and $T_{1}$ is the partial transposition on the first system on a given local basis.
As mentioned before, using the diamond norm, we can replace the hard equation in the SDP for P-Divisibility with a smoother yet physically meaningful alternative, similar to what we did for the SDP determining the CP-Divisibility. This is the optimization problem we are using in the next examples, and it is formulated as follows:
P-Divisibility with Diamond Norm (Alternative for Qubits)
$\text{Given}\leavevmode\nobreak\ \varrho_{2|0},\,\varrho_{1|0}$ $\operatorname*{Minimise}_{\varrho_{2|1},\,\sigma_{A},\,\sigma_{B}}\|\mathcal{J}^{-1}(\varrho_{2|0})-\mathcal{J}^{-1}(\varrho_{2|1})\circ\mathcal{J}^{-1}(\varrho_{1|0})\|_{\diamond}$ s.t. (21) $\qquad\varrho_{2|1}=\sigma_{A}+\sigma_{B}$ $\qquad\sigma_{A}\geq 0$ $\qquad\sigma_{B}\mbox{ is the Jamio\l kowski image of }\Lambda_{2|1}\circ T_{1}$ $\qquad\mbox{Tr}_{\mathcal{H}_{2}}\left[\varrho_{2|1}\right]=\mathds{1}_{\mathcal{H}_{1}},$ where $\Lambda_{2|1}$ is CPTP, $T_{1}$ is the partial transposition on the first system on a given local basis and $\mathcal{J}^{-1}(\varrho_{k|0})$ is the map associated with $\varrho_{k|0}$ via the (Choi)-Jamiołkowski isomorphism.
@@ V Witnessing non-Markovianity in a Quantum Dynamics
Non-markovian effects are commonly associated with a backflow of information [13]. One of the most typical measures of it relies on the increase of distinguishability of two states [15]. Differently, the formalism proposed here does not rely on any witness of the backflow of information and does not request any restriction on the dynamics. To show its relevance, we will apply it to three paradigmatic models.
The first example is a collisional model. This class of model has been initially proposed as a toy model to simulate any Markovian dynamics. Recently, by allowing the generation of correlations in the environment, collisional models have been proven to be a powerful tool in the studies of non-Markovian dynamics (for a review, see [32]). In this model the environment is described by separated particles and the system interacts with each particle at once and sequentially—see Fig. 2. The advantage of this model is that the three main ingredients that may lead to a non-Markovian evolution, namely the correlations in the environment, its internal dynamics, and the detailed interaction between the system and the environment can be separately fully controllable. Indeed, for qubits, collisional models can simulate any non-Markovian dynamics [51].
@@
The third example is a generalization of the second model. A random unitary evolution or a convex combination of unitary channels. This is a broad family of channels with a handy representation for higher dimensions [52].
@@ V.1 Collisional Model
Two main ingredients are common in all collisional models: time is discretized, and the environment is considered to be formed by elementary subsystems. Each encounter between the system and the $j$-th particle in the environment is denoted by the unitary operator $U_{j}$, acting for a fixed duration $\tau$. The environment’s particles are typically identical and exhibit no interactions. The resulting evolution after $j$ collisions is expressed as $\rho_{j\tau|0}=\text{Tr}_{j}...\text{Tr}_{1}[U_{j}...U_{1}(\rho\otimes\omega_{env})U_{1}^{\dagger}...U_{j}^{\dagger}]$. A pictorial representation of this process is displayed in Fig. 2. This process simulates a Markovian dynamics. However, for non-Markovian dynamics, as indicated in [51, 53, 54], the environment’s particles can no longer remain noninteracting; instead, they must exhibit some level of correlation.
@@
An advantage of using SDPs to quantify P- and CP-non-divisibility is that the best channel approximating the optimal distance is obtained as a by-product. In some cases, however, the resulting best channel has been observed to be the identity, meaning that no other channel will improve in approximating $\Lambda_{2|0}$ starting at $\Lambda_{1|0}$, $\Lambda_{1|0}$ being already the closest channel as a factor composing the former. This induces a notion of absolute non-divisibility and a possible justification for the result is introduced in Appendix D.
@@
Figure 5: Distance to divisibility for the dephasing channel. Dynamics starts at time $t_{1}$ and ends at time $t_{2}$ (with values in the range $0,2\pi]$ and $t_{2}\geq t_{1}$). On the left panel, divisibility for a selection of initial times is shown. Notice that, as $t_{1}$ increases, the sooner non-divisibility starts, with extended regions of exact divisibility occurring only for $t_{1}\leq\pi$. Compared with figure 4, it can be seen that divisibility can only be attained when $p(t_{2})>p(t_{1})$.
@@ V.2 Dephasing Dynamics
The dephasing channel is a completely quantum noise channel, where quantum information is lost without the loss of energy. It models fiber-optic communication channels [56] and superconducting circuits [57], impacting quantum communication and quantum computation.
@@
with $p(t)=e^{4}\,(1-e^{-2(1-\cos(t))})/(e^{4}-1)$ (see Fig. 4). The corresponding master equation is $\dot{\rho}=\frac{\gamma(t)}{2}(Z\rho Z-\rho)$ with $\gamma(t)\leq 0$ for $t\geq\pi$ and consequently non-Markovian [52]. Another interesting feature of this model is that for $p=1$ the map is non-invertible; this is a regime where many different Markovian witnesses are not well defined [19, 20, 21]. The time parameter here is again continuous, but note that we will always be interested in three different time steps: the initial time $t_{0}$, the intermediate time $t_{1}$, and the final time $t_{2}$.
@@
In systems with dimension higher than two, one possible way to generalize the dephasing channel is given by replacing the random action of the Pauli $Z$ unitary with projections on the elements of the computational basis, $|\Pi_{k}=\ket{k}\bra{k}$ [58]. The effect of the dephasing channel on a density matrix $\rho$ is given by
$\Lambda^{\textrm{dep}}_{p}(\rho)=(1-p)\rho+p\,\sum_{k}\Pi_{k}\,\rho\,\Pi_{k},$ (24)
where the channel is now directly parametrized by the probability of getting a projection, $p\in[0,\,1]$.
Notably, also here the composition of two dephasing channels return a third dephasing channel, with dephasing probability greater than the maximum of the two input probabilities. Direct calculation shows that $\Lambda_{p^{\prime}}\circ\Lambda_{p}=\Lambda_{p+p^{\prime}-pp^{\prime}}$, and $p+p^{\prime}-pp^{\prime}\geq\max\{p,p^{\prime}\}$. This implies in particular that, for $\Lambda_{1|0}=\Lambda^{\textrm{dep}}_{p}$ and $\Lambda_{2|0}=\Lambda^{\textrm{dep}}_{q}$, a solution to the divisibility problem is always obtainable when $q\geq p$, by using $p^{\prime}=(q-p)/(1-p)$, for $p<1$, and setting $p^{\prime}=0$, for $p=1$. Indeed, in running the program of Eq. (IV) for channels $\Lambda_{1|0}=\Lambda^{\textrm{dep}}_{p}$ and $\Lambda_{2|0}=\Lambda^{\textrm{dep}}_{q}$ in input dimensions $3$ and $5$, the maps obtained as a result of the optimization coincide with the analytical expectation. In the cases where CP-divisibility is not achieved, namely for $q<p$, the identity channel is obtained, indicating a strong character of non-invertibility from channels with higher values for the dephasing parameter to lower values. Results for CP-nondivisibility for dimension $5$ are shown in Fig. 6, where it can be observed both that $\|\Lambda_{2|0}-\Lambda_{2|1}\circ\Lambda_{1|0}\|_{\diamond}=\|\Lambda_{2|0}-\mathcal{I}\circ\Lambda_{1|0}\|_{\diamond}=(q-p)\,\|\mathcal{I}-\sum_{k}\Pi_{k}(\cdot)\Pi_{k}\|_{\diamond}$ for $q<p$, and equal to zero, within numerical precision, for $q\geq p$.
Figure 6: CP-nondivisibility for different values of parameters $p$ and $q$ in a space with dimension 55. The onset of divisibility occurs exactly at $q=p$, with the resulting intermediate map confirmed numerically to be $\Lambda_{2|1}=\Lambda^{\textrm{dep}}_{(q-p)/(1-p)}$.
V.3 Convex combinations of unitary channels
Moving away from physically motivated channels, in this section we consider a class of channels that is more interesting from a mathematical perspective to provide an analysis of scalability of the method. Namely, we consider the particular subset of unital maps formed as convex combinations of unitary channels. It is a relevant and broad family of channels that encompasses also physically relevant maps, the dephasing channels of the previous subsection being a notable example thereof [44].
In full generality, the maps are defined as
$\Lambda_{\bm{p},\,\bm{U}}(\rho)=\sum_{i=1}^{n}p_{i}\,U_{i}\,\rho\,U^{\dagger}_{i},$ (25)
for a set of unitaries $\bm{U}=\{U_{1},\ldots,U_{n}\}$ and a given probability distribution $\bm{p}$. In the particular cases considered here, we analyse the situation where $\Lambda_{2|0}=\mathcal{I}$ and $\Lambda_{1|0}$ is a randomly generated channel with unbiased combinations of the unitaries: $\bm{p}=\{1/n,\ldots,1/n\}$, and $U_{i}$ are drawn using the Haar measure.
The SDP is run in scenarios with differing values of the parameters. We considered Hilbert spaces of dimensions $d=2,\ldots,7$ and, for dimensions $2$ and $3$, we analyse the impact of using different number of unitaries that conform the channel, $n$. For each particular selection of parameters, $500$ channels $\Lambda_{1|0}$ are sampled and the corresponding CP-nondivisibility is computed. The optimization for each channel is timed, average values obtained for $n=5$ and for different values of the dimension are shown in Fig. 7. The values are compatible with a polynomial scaling, with a practical scaling of order $d^{5}$, as observed in the agreement with the curve fitted over the computed data, shown in the figure.
Figure 7: Average time taken to run the SDP for CP-nondivisibility for each channel $\Lambda_{1|0}$ sampled. A behavior scaling with $d^{5}$ is observed in practice, as shown by the fitted polynomial over the computed data.
With respect to changes in $n$, no noticeable change is observed, as shown in table 1, which reveals that, for typical cases, particular instances of the channels used do not impact the computation of the SDP. All computations were performed on an Intel Xeon W-2295 workstation operating at $3.0$ GHz, using the MOSEK optimizer [59].
$d$ $n$ 11 22 55 1515
22 $0.11\pm 0.03$ $0.10\pm 0.02$ $0.10\pm 0.03$ $0.10\pm 0.03$
33 $0\pm 0.06 0.33±0$$3\pm 0.07$ $0.29\pm 0.06$ $0.32\pm 0.06$
Table 1: Average time taken to solve each SDP in seconds, for different values of $n$ and dimensions d=2, 3d=2,\,3. Values are comparable for fixed dimensions, showing that the technique is not affected by particular instances of input channels used.